OpenStudy (anonymous):

Which is the simplified form of 18 times a to the sixth power all over 12 times a squared.? 6a3 6a4 3 over 2. a3 3 over 2. a4

OpenStudy (anonymous):

18a^6/12a^2 simplify the fraction a^6/a^2=a^4 18/12=3/2* 18a^6/12a^2=3/2*a^4